A projectile launch speed is five times its speed at the maximum height,find the launch angle teta?
1
Expert's answer
2019-12-18T11:32:44-0500

Let us assume that the launch angle be "\theta"  and projectile launch speed be "v"

then the horizontal component of velocity = "vcos\theta"

and the vertical component of velocity = "vsin\theta"

then the speed of the object at maximum height = "vcos\theta"

It is given that projectile launch speed is five times its speed at the maximum height

⟹ "v=5vcos\theta"

⟹ "cos\theta=1/5"

"\theta=78.5" 0